### v2

# Astrophysical signals from the neighboring brane

Abstract
In this paper we consider the class of extra-dimensional models with two branes and show that each field have to be located on both neighboring branes, thus forming two independent effective fields with different masses. Specific results are illustrated in the framework of a previously developed model. We show that the fermion masses on the two branes are different at the low energies because the fine-tuning of the parameters is necessary only on the observer-populated brane (brane-1). The second brane (brane-2) has no observers and hence the fine-tuning is not obligatory leading to the particle masses of the order of the initial energy scale of the universe formation. The gauge fields are uniformly distributed in the bulk, ensuring charge universality. The Higgs mechanism of masses formation on the both branes is discussed.
Stable, massive charged leptons on brane-2 can be considered as components of the dark matter due to their significantly reduced Compton scattering cross section. Furthermore, we show that inter-brane interaction via photons and gravitons allows the massive fermions on the second brane to be a source of ultra-high energy particles.
11footnotetext: e-mail: sergeirubin@list.ru
1 Introduction
The idea of multidimensional gravity is a vital tool for obtaining new theoretical results and explaining known phenomena ([Abbott:1984ba, Brown:2013fba, Bronnikov:2009zza, Chaichian:2000az]). Over the course of several decades, numerous significant results have been achieved, and various concepts have been developed. The hierarchy problem is undoubtedly one of the most significant issues (see e.g. [Gogberashvili:1998vx, 1999PhRvL..83.3370R, ArkaniHamed:1998rs]). The problem of the small cosmological constant is discussed in a variety of papers, see e.g. [Krause:2000uj, Bronnikov:2023lej], while the concept of multidimensional inflation is discussed in detail in several sources, including [2002PhRvD..65j5022G, Bronnikov:2009ai, Fabris:2019ecx]. These works assume the extra-dimensional metric is static and stabilization occurs on high-energy scales. The stabilization of extra space as a purely gravitational effect has also been studied in [2002PhRvD..66d4014G, 2003PhRvD..68d4010G, Arbuzov:2021yai].
[…]
For several decades, significant interest has been observed in theories concerning the localization of fields on a three-dimensional hypersurface, known as a brane, within a multidimensional spacetime. Akama was the first who introduced this concept [Akama:1982jy]. The hypothesis of the thick brane was independently formulated by Rubakov and Shaposhnikov [Rubakov:1983bb].
Thin branes, widely discussed in the literature, suffer from some intrinsic problems [Burgess:2015kda]. Therefore, despite hopeful results, they are gradually being replaced by branes with internal structure - thick branes [Bronnikov:2006bu, Chumbes:2011zt, Hashemi_2018, Dzhunushaliev:2019wvv, Bazeia:2022vac, Wan:2020smy]. The latter require special conditions to prove their existence and stability. This could be the choice of a special metric, including the warp factor [Oda_2000] or a special form of the scalar field potential [Bazeia:2022vac]. A purely geometrical approach based on $f(R)$ gravity in 5D is given in [Guo:2023mki], where a considerable number of bibliographic references can be found.
In brane-world scenarios, the key issue is the localization of fields on branes and the recovery of effective 4D gravity. The localization of gauge and spinor fields on branes is a necessary element of any brane model and is a widely discussed topic in the literature. The solutions representing the tensor mode excitations in the presence of a brane are examined in [Cui:2020fiz, Xu:2022xxd] with a focus on the stability problem.
The introduction of the warp factor in a D-dim metric [Oda_2000] generalizes the analysis, especially in the fermion sector, as outlined in [Wan:2023usr], which provides a detailed analysis of this subject in even-dimensional space-time. Fermions in other types of extra metrics are studied in [Gogberashvili_2007, Dantas:2015dca].
The two-brane ideology is often based on the well-known 5D Randall-Sundrum model (RS1), describing two 3D thin branes located at two fixed points of the orbifold. The model postulates a fixed distance between the branes, which is one of its disadvantages. A substantial amount of references can be found in [Olechowski:2024wcf, Tanaka:2000er, Feranie_2010]. The formation of multiple thick branes is considered in, e.g. [Bronnikov:2007kw, Dzhunushaliev:2019wvv]. In this study, we base on the results of our paper [Popov:2024nax] where a new class of branes was discussed. Our analysis there shows that one-brane metrics are exceptions rather than the norm, and we are particularly interested in the set of metrics that describe two-brane configurations. It is shown here that both branes are filled by the same sorts of fields because they have a common origin at the highest energies. In the course of energy reduction, a field is split into two effective fields with different masses depending on the branes properties. Particles located on different branes can interact by exchanging photons. This opens up the possibility of clarifying phenomena such as dark matter [Song:2023xdk] and cosmic rays of the highest energies. Effective superluminal propagation of massless signals in 4D is studied in [Dai:2023zsx].
As was shown in [Bronnikov:2023lej], the Sandwich model of extra dimensions discussed here could serve as the basis for resolving the hierarchy problem, including the issue of the smallness of the cosmological constant.
The paper is structured as follows. In Section 2, we discuss the energy scales suitable for the formation of classical extra-dimensional metrics and the particle content on branes. Section 3 introduces the basis of the Sandwich model of extra dimensions. Section 4 discusses fermions, photons, and the Higgs field distribution over the extra dimensions. We show that the spinor fields split between the two branes. It contains the analysis in which cases the matter located on different branes can be described as independent fields. Section 5 discusses the observable signals from the uninhabited brane and the Higgs mechanism of the mass generation on the both branes. Finally, we give a Conclusion in Section 6.
2 Spreading of matter on branes in the course of their formation
2.1 Scale of the branes formation
The widely accepted inflationary paradigm implies that the formation of the physical laws takes place above the inflationary scale $H_{I}\sim 10^{14}$eV. Therefore, action describing field dynamics appears at extremely high energies where the quantum fluctuations fade out. The Rundall-Sundrum model with two branes (RS1 model) and its developments [Randall:1999vf] use the second brane as the source of gravity. But there is an aspect which is usually omitted. Indeed, the static two-brane structure does not exist forever, but was formed at high energies. When the universe cools down, two processes occur simultaneously: the extra-dimensional metric describing the branes stabilizes and the fields are randomly settled on them. As a result, the matter fields must be present on all branes.
The two-brane structure is broken by external perturbations if the energy scale is higher than the inverse distance $l$ between the two branes. So we can consider static metrics at the energies where the wavelengths of the 4D fluctuations are much larger than the extra-space size $l_{extra}$. The characteristic wavelengths at the de Sitter stage are of the order of the inverse Hubble parameter $H^{-1}$. Therefore, the estimation $H\equiv H_{D}\ll l_{extra}^{-1}$ provides us with the energy scales where the specific two-brane metric can be considered static.
The D-dim Planck mass $m_{D}$ determines the highest possible energy scale. So the first limit is $l_{extra}^{-1}\ll m_{D}$. Another inequality can be found by using the inflationary concept. It was shown in [Nikulin2019] that a scale of compact extra dimensions cannot be larger than the inflationary scale, $l_{extra}\ll 1/H_{I}$, in order not to destroy the slow rolling during the inflation. Combining these inequalities, we get the interval for the extra-space scale variation
$H_{I}\ll l_{extra}^{-1}\ll m_{D}$ (1)
The whole picture looks as follows. The extra-dimensional metric is formed below the energy scale $m_{D}\sim 10^{18}$eV, and is stabilized above the scale $H_{I}\sim 10^{14}$eV. It remains static at lower energies. The initial fluctuations of the fields are responsible for the stochastic distribution of matter on both branes.
Space-time was created at sub-Planckian energies, as is commonly assumed. This time period is characterized by strong fluctuations of the metric and the fields. The static extra-dimensional metrics, such as the two-brane structure, are formed at much lower energies and are accompanied by fields settling on the branes. The physical parameters characterizing the behavior of a field are contingent upon the properties of the branes.
The hierarchy of the energy scales and the following fine-tuning are necessary phenomena for the formation of complex structures [2007unmu.book..231D]. The selection of the physical parameters for the fine-tuning on a brane is a rather delicate process. It can be done only for one brane and it is reasonable to choose the brane-1 which is populated by observers. Brane-2 is then populated by all known sorts of particles with masses much heavier than those on brane-1. As discussed above, the physics as a whole and the brane structure in particular are formed at high energy scale in the interval $10^{14}-10^{18}$eV. This means that the natural masses of particles on the brane-2 are of the same order of magnitude.
We come to conclusion that there are all sorts of known particles on the brane-2 which are many order of magnitude more massive compared with those on the brane-1. The matter on the brane-2 is incapable of forming complex structures like stars because of the lack of the fine-tuning.
The potential contribution of such massive particles to the dark matter sector warrants investigation. The following subsection presents a general analysis of this hypothesis, decoupled from specific brane-world configurations. We assume that the gravitational and electromagnetic forces between matter are invariant with respect to the brane on which the matter is located. This assertion, was rigorously established in [Popov:2024nax].
2.2 Dark matter on the uninhabited brane
One of the main goals of the cosmology is to understand the nature of the dark matter. Special efforts have been applied to explain the properties of this phenomenon and the origin of dark matter [Gani:2014lka, 2006PhRvD..73a5001B, 2007PhR...453...29H, 2006PZETF...83..3, Arbuzova:2021etq, Belotsky:2017pcf, Belotsky:2014kca, 2003PhRvD..68j3003B]. Many theoretical models as well as the observational evidences can be found in the reviews [Beylin:2019gtw, Arbey:2021gdg, Misiaszek:2023sxe]. The idea of extra dimensions is also attracted in the framework of this discussion, [Chu:2011be, 2005JCAP...07..001B, 2009MPLA...24..667K, 2006PhRvD..73a5001B]. Some of them are treated the brane idea [Boehmer:2007xh, SHAHIDI_2011] which is natural basis for production of weakly interacted particles. The concept of the mirror world [Beradze:2019yyp, Dubrovich:2021gdn] can be attracted for the same reason.
The most probable content of heavy particles on the brane-2 is the leptons, protons, neutrinos, and their antiparticles. Among these particles, heavy neutrinos emerge as promising candidates for the dark matter component.
Let us return to the two-brane structure discussed above. We have shown that the observers could be on the brane-1, while the heavy electrons with the standard charge are on the brane-2.
A more thorough discussion is necessary to address the implications of charged particles, such as electrons and protons. The central question pertains to the potential role of these charged heavy electrons and protons in the context of dark matter. The primary challenge lies in understanding their interaction with photons, see [Kamada2016EffectsOE]. Such particles, being extremely massive are non-relativistic and could form bonded states like positronium and the hydrogen atom. The positronium life-time $\sim m^{-1}$, ($m$ is the heavy electron mass on the brane-2) and hence decay instantly after its formation. The concentration of the heavy electrons is substantially decreased, yet it remains nonzero. So we have to estimate the Compton effect on brane-2, i.e. the photon scattering on the non-relativistic heavy electron.
The cross section of the Compton effect depends on a particle mass $m$ as
$\sigma\sim\frac{1}{m^{2}}\sim 10^{-26}\text{cm}^{2}.$ (2)
The second approximate equality is the known experimental result for the standard electron mass $m_{e}$. The same cross section for the heavy electrons is smaller by the factor $(m_{e}/m_{2})^{2}$ than the standard one. Here $m_{e}$ and $m_{2}$ are the masses of the ’’normal’’ electron and the ’’heavy’’ electrons. We can estimate the mass of the heavy electrons supposing that their Compton cross section should be smaller than the lower observational bound $\sim 10^{-44}$ cm2 for the dark matter cross section,
$\sigma_{2}\sim\frac{1}{m_{2}^{2}}\sim\frac{m_{e}^{2}}{m_{2}^{2}}\sigma_{e}\lesssim 10^{-44}\text{cm}^{2}.$ (3)
It is easy to see that the heavy electron mass should be really large,
$m_{2}>10^{9}m_{e}\sim 10^{6}\text{GeV},$ (4)
where the value of $\sigma_{e}$ is taken from (2). The mass estimate (4) for heavy electrons on brane-2 allows one to consider them as the dark matter candidate. There are several possible candidates for the role of dark matter: hydrogen, neutrinos, electrons and protons. All of them are located on the brane-2, and hence are very heavy.
The results of this section are based on several assumptions. a) The two-brane metric is the static solution of the classical equations. b) The gravitational and electromagnetic interactions are the same on both branes and uniformly distributed between them. The model elaborated in [Popov:2024nax] is briefly presented below to support points a) and b).
c) Each matter field is necessarily located on two branes and is observed as two independent fields with different masses. This point is discussed in the subsections 4.2 and 4.3.
3 Two-brane metric. Sandwich model of extra dimensions
3.1 The model choice
The presence of at least two branes is necessary element of this study so that we will base on the model elaborated in [Popov:2024nax]. In that study, we impose the extra-dimensional metric and show that a two-brane structure is formed under some assumptions which are not very restrictive. The metric is supposed in the form
$ds^{2}=\text{e}^{2\gamma(u)}\left[dt^{2}-\text{e}^{2Ht}(dx^{2}+dy^{2}+dz^{2})\right]-du^{2}-r(u)^{2}d\Omega_{n-1}^{2},$ (5)
where $d\Omega_{n-1}^{2}$ is the metric on a unit $n-1$-dimensional sphere. The radial extra coordinate $X^{4}\equiv u$ and $X^{\mu}\equiv x^{\mu},\mu=0,1,2,3.$ The angular coordinate $X^{5}=\theta$ varies in the finite region. Here we consider a 6D space so that $D=4+n,\,n=2$ and the set of extra-dimensional coordinates is $y=\{X^{4},X^{5}\}\equiv\{u,\theta\}$. We do not fix the limits of the radial coordinate $u$ a priori. The Hubble parameter $H$ is arbitrary.
Consider $f(R)$ gravity in a $\text{D}=4+n$-dimensional manifold $M_{\text{D}}$:
$S=\frac{m_{\text{D}}^{\text{D}-2}}{2}\int_{M_{\text{D}}}d^{\text{D}}X\sqrt{|g_{\text{D}}|}f(R),\quad f(R)=aR^{2}+R+c,$ (6)
where $g_{\text{D}}\equiv\mathop{\rm det}\nolimits g_{\text{M}\text{N}}$, $\text{M},\text{N}=\overline{1,\text{D}}$, the $n-1$-dimensional manifold $M_{n}$ is assumed to be closed, $f(R)$ is a function of the D-dimensional Ricci scalar $R$, and $m_{\text{D}}$ is the D-dimensional Planck mass.
Remark that there are still no traces of branes, neither in the metric (5) nor in the gravity action (6). Nevertheless, branes are inevitably formed during the cooling of the universe. See [Popov:2024nax]. The two-brane structure reveals as solutions of the classical equations
$-\frac{1}{2}{f}(R)\delta_{\text{N}}^{\text{M}}+\Bigl{(}R_{\text{N}}^{\text{M}}+\nabla^{\text{M}}\nabla_{\text{N}}-\delta_{\text{N}}^{\text{M}}\Box_{\text{D}}\Bigr{)}{f}_{R}=0$ (7)
with $f_{R}={df(R)}/{dR}$, $\Box_{\text{D}}=\nabla^{\text{M}}\nabla_{\text{M}}$.
The static extra-dimensional metrics, being solutions of Eq.7, form an infinite set, as discussed in the papers [Rubin:2014ffa, Rubin:2015pqa, Nikulin2019]. It was shown there that their common form is characterized by two zeros of the metric functions $r(u_{1,2}),\gamma(u_{1,2})$, which is illustrated in Figs.1. The manifold is a compact n-dimensional space of a finite volume.
Figure 1: The typical metric functions, $\gamma(u)$ solid line, $r(u)$ dashed line. The interval is normalized so that the endpoints are $u_{1}=0,\quad u_{2}=\pi$ in the units $m_{D}=1$. Their shapes depends on additional conditions and are characterized by zeros at the endpoints.
The classical equations are senseless at the distances $m_{D}^{-1}$ where the quantum regime dominates. This statement holds for any classical system and is the reason why the space-time quantum foam exists at the Planck energy scale. In our case, this means that a brane width should be substantially larger than $m_{D}^{-1}$.
The thickness of the brane is related to the fields localization in its vicinity. As was shown in [Bronnikov:2023lej, Popov:2024nax], the classical field distributions along the extra coordinate $u$ are characterized by two areas of localization. One can conclude that the model described above leads to the formation of the two-brane structure and we will refer to it as the Sandwich Model of Extra Dimensions (SMED).
3.2 The Planck mass
In this paper, we widely use the units $m_{D}=1$ and it is instructive to restore its dimensionality at the final stage. The $D$-dimensional Planck mass $m_{D}$ can be fixed by its connection to the known four-dimensional Planck mass $m_{4}$. For this purpose, we make the necessary preparations here, starting with the representation of the Ricci scalar of 6D extra space in the form
$R\equiv R_{6}=e^{-2\gamma(u)}R_{4}(x)+R_{2}(u)+\mathcal{R}(u),$ (8)
$\mathcal{R}=-\frac{8}{r(u)}\gamma^{\prime}(u)r^{\prime}(u),$ (9)
$R_{2}(u)=-\frac{2r(u)^{\prime\prime}}{r(u)},$ (10)
where 2D extra metric presented in (5) has been taken into account. Here $R_{4}$ is the Ricci scalar of our 4D space and $R_{2}$ is the Ricci scalar of the extra space. We also assume that the distributions of matter and metric functions over the extra dimensions depend only on the radial extra coordinate $u$. In this case, the action (6) takes the form
$S=\frac{m_{\text{D}}^{\text{D}-2}}{2}\int_{M_{\text{D}}}d^{\text{D}}X\sqrt{|g_{\text{D}}|}\,f(e^{-2\gamma(u)}R_{4}(x)+R_{2}(u)+\mathcal{R}(u))\simeq$
$\frac{m_{\text{D}}^{\text{D}-2}}{2}\int_{M_{\text{D}}}d^{4}xd^{n}y\sqrt{|g_{\text{D}}|}\,[f(R_{2}(u)+\mathcal{R}(u))+e^{-2\gamma(u)}f_{R}(R_{2}(u)+\mathcal{R}(u))R_{4}(x)+O(R_{4}^{2})]\simeq$
$\frac{m_{4}^{2}}{2}\int d^{4}x\sqrt{|g_{4}|}[R_{4}(x)-2\Lambda]$ (11)
The 4D Planck mass $m_{4}$ is related to the D-dim Planck mass $m_{D}$ as follows:
[…]
where $v_{n-1}$ is the volume of an $n-1$-dim sphere of unit radius
$v_{n-1}=\frac{\pi^{\frac{n-1}{2}}}{\Gamma(\frac{n-1}{2}+1)}.$ (13)
The Planck mass $m_{4}$ is measured at the low energies, and we assume that the inequality $R_{4}\ll R_{2}$ holds.
Let us estimate the value $m_{D}$. To this end, suppose that all dimensional values are of the order of $m^{D}$, so that
$m_{D}\sim m_{4}/\sqrt{v_{n-1}}.$ (14)
The volume (13) has maximum $v_{n-1}\simeq 2500$ at $n\simeq 20$ which gives the estimation of the smallest value of $m_{D}\sim 0.02m_{4}\simeq 10^{17}$eV. Here and below we suppose $m_{D}\sim 10^{17}\div 10^{18}$eV and the scale of the two-brane structure $l_{extra}\sim 10^{-14}\div 10^{-15}$eV-1 for the estimations. Other terms in (3.2) contribute to the vacuum energy density $\Lambda$, see discussion in [Bronnikov:2023lej] , which is not used here. The knowledge of the metric functions $\gamma(u),r(u)$ as in Figure 1 permits to restore the 4D Planck mass $m_{D}$, see expression (14).
A non-trivial scenario arises if the number $n$ of extra dimensions is large. In this case, the volume of extra dimensions (13) is a rapidly decreasing function of the number of extra dimensions. For example, $v_{n}\simeq 4\cdot 10^{-15}$ at $n=100$, so
$m_{D}\sim 10^{8}m_{4}\sim 10^{27}\text{GeV}$
The onset of the quantum regime occurs at a scale $1/m_{D}\sim 10^{-8}/m_{4}$, which is significantly smaller than the typically mentioned estimation $1/m_{4}$.
4 Fields splitting between two branes
Here we study the distributions of fields over the two branes and their features leading to observational consequences after the reduction to the 4D physics. We also discuss conditions under which a specific field disposed on two branes can be considered as two fields with effectively different parameters.
The reduction to 4 dimensions assumes integration over the extra coordinates. So we have to know the distributions of fields (scalars, fermions, vectors) within the extra dimensions. The suitable way of a field decomposition is as follows:
$\Phi(x,y)=\Upsilon(x)Y_{cl}(y)+\sum_{k=0}^{\infty}\Upsilon_{k}(x)Y_{k}(y)\simeq\Upsilon(x)Y_{cl}(y),$ (15)
where $\{Y_{k}(y)\}$ is a set of orthonormalized functions. It is usually assumed that they satisfy the equation
$\Box_{n}Y_{k}(y)=\lambda_{k}Y_{k}(y),k=0,1,2,...$ (16)
The zero mode $Y_{0}(y)=1$ satisfies this equation with $\lambda_{0}=0$. The function $Y_{cl}(y)$ is the classical solution to the field equation(s) provided that the derivatives in $x$ coordinates of our space are neglected. This approximation holds if the wavelengths of the perturbations acting in our 3-dim space are much larger than a scale of the extra dimensions. It also means that the sum describing the field fluctuations along the extra dimensions can be omitted. The function $\Upsilon$ is assumed to be endowed with group indices depending on the group representation of the field while the scalar function $Y_{cl}$ is responsible for the distribution over the radial extra dimension. The fields are treated as trial functions in this research. This step does not involve any fine-tuning. The amplitude of the classical distributions, $Y_{cl}(u)$, is of the order of the D-dimensional Planck scale, $m_{D}$.
4.1 Fermions distribution over the extra dimensions. Analytical solution
In this subsection we shortly remind the results of [Popov:2024nax] concerning the distribution of fermions along the radial extra-dimensional coordinate $u$ near the brane and in the bulk. General form of the free massless fermion action is
$S_{fermion}=\int d^{D}z\sqrt{|g_{D}|}\bar{\Psi}i\Gamma^{M}D_{M}\Psi$ (17)
The mass term is assumed to be generated by the interaction with the Higgs field. The particular form of the curved gamma matrices $\Gamma^{M}$ and the covariant derivatives can be found in [Oda_2000]. The useful discussion is also given in [Appelquist:2000nn, Randjbar-Daemi:2000lem]. The more thorough model in even dimensions is outlined in [Wan:2023usr]. That paper studies conditions for decoupling the left and right chiralities of spinors. A useful discussion is also given in [Appelquist:2000nn, Randjbar-Daemi:2000lem]. A 6D spinor
$\Psi=\left(\begin{array}[]{c}\xi_{1}\\ \xi_{2}\end{array}\right)$ (18)
acting in six dimensions is equivalent to a pair of 4-dimensional Dirac spinors $\xi_{1}$ and $\xi_{2}$. As shown in [Gogberashvili_2007], both 4D spinors, $\xi_{1}$ and $\xi_{2}$, are equal in the s-wave approximation used in our approach. In this paper, we are interested in the spinor distribution over the extra dimensions. The spinor $\psi$ is characterized by both spinor and group indices and functions within a six-dimensional framework. The field decomposition (15) in this case reads
$\Psi(x,y)=\psi(x)Y_{f}(u).$ (19)
Here the normalized function $\psi(x)$ is endowed with spinor and group indices acting in 4-dimensional space.
This study focuses on the distribution of fields across the extra-dimensional coordinates rather than their reduction to four-dimensional physics. Consequently, our attention is directed towards the form of the function $Y_{f}(u)$ which is a solution to equation
$\Big{[}\partial_{u}+\frac{r^{\prime}(u)}{2r(u)}+2\gamma^{\prime}(u)\Big{]}Y_{f}(u)=0.$ (20)
The analytical solution to this equation has the form
$Y_{f}(u)=c_{2}r(u)^{-1/2}e^{-2\gamma(u)}.$ (21)
Figure 2: Fermion distribution over the extra dimensions. The metric functions are represented in Fig.1. The localizations near $u_{1}=0$ and $u_{2}=\pi$ are evident.
The particular distribution over the radial coordinate $u$ is shown in Fig.2. The more detailed analysis is presented in [Popov:2024nax]. The distribution has the sandwich-type structure with sharp peaks at the coordinates $u_{1}$ and $u_{2}$ and almost zero field value between them. The region proximate to the coordinate $u_{1}$ is designated brane-1, and the area near the coordinate $u_{2}$ is labeled brane-2.
4.2 Spreading a field over two branes and splitting it into two independent effective fields.
Consider an action featuring the specific field $\zeta(x,y)$ along with its associated set of Lagrangian parameters $\{\lambda_{in}\}$, which include mass and coupling constants. Let us find the conditions for which this field is observed at low energy as two independent fields characterized by different sets of parameters $\{\lambda_{in}\}\to\{\lambda_{1,2}\}$. The transition amplitude is calculated using the path integral approach,
$A=N\int D\zeta(x,y)\exp{\left(iS[\zeta;\lambda_{in}]\right)},\quad D\zeta(x,y)=\prod_{x,y}d\zeta(x,y).$ (22)
Here $\lambda_{in}$ is an initial set of physical parameters. An action $S$ in (22) is a functional of a function $\zeta$ acting in D-dim. At the highest energy levels, where $H\sim m_{D}$, this field is dispersed throughout the entire extra space. However, as the system’s energy decreases to lower levels, the scenario changes. As was discussed in the subsection 2.1, the extra space structure is formed at the scale $l_{e}\lesssim 10^{-27}$cm. At the present time, the smallest scale where achieved at the colliders $l\sim 10^{18}$cm. The inequality $l_{e}\ll l_{s}$ allows to neglect the 4D variation of fields and use the classical equations for fixing the extra-dimensional metric and the fields distribution over extra coordinates $y$. The decomposition (15) in the form
$\zeta(x,y)=\Upsilon(x)Y(y)$ (23)
appears to be useful here. So we can apply the stationary phase method and substitute the distribution $Y(y)$ by the classical one $Y(y)\to Y_{cl}(y)$ over the extra dimensions and leaving the 4D distributions $\Upsilon(x)$ arbitrary. The example of the classical equation admitting analytical solution due to its simplicity, is written in the subsection 4.1. The amplitude has the form
$A=N\int D\zeta(x,y)\exp{\left(iS[\zeta;\lambda_{in}]\right)}=N\int\prod_{x}d\Upsilon(x)\prod_{y}dY(y)\exp{\left(iS[\Upsilon(x)Y(y);\lambda_{in}]\right)}\simeq$
$N^{\prime}\int\prod_{x}d\Upsilon(x)\exp{\left(iS[\Upsilon(x)Y_{cl}(y);\lambda_{in}]\right)}=N^{\prime}\int D\Upsilon\exp{\left(iS^{(4)}[\Upsilon(x));\lambda_{eff}]\right)},$ (24)
where the main order of the approximation is used in the last line. Knowledge of the function $Y_{cl}(y)$ permits to integrate over the extra dimensional coordinate $y$ which lead to the renormalization of the initial physical parameters values $\lambda_{in}\to\lambda_{eff}$. $S^{(4)}$ in effective 4D action. The general conclusion is that the low energy effective parameters $\lambda_{eff}$ appears to be dependent on accidental field distribution at high energies. This result being very important, has been performed for various fields in [Rubin:2016ude, Petriakova:2023klf, Bronnikov:2023lej, Popov:2024nax].
In the papers mentioned above, it was also shown that the spinor, scalar and the Higgs field distributions has sharp peak near end points, while the intermediate region is characterized by extremely small field value. Fig.2 illustrates the spinor field localization near each endpoint. This remark opens new opportunity in the research and we consider it more thoroughly below. The region between end points $(u_{1},u_{2})$ can be divided into three distinct parts. The first region $(u_{1},u_{1}^{\prime})$, denoted as $\mathcal{U}_{1}$, encompasses the extra-dimensional radial coordinate $u\in\mathcal{U}_{1}$ forming the brane-1, while another region $(u_{2}^{\prime},u_{2})$ includes the coordinate $u\in\mathcal{U}_{2}$ and forms brane-2. Both of this regions are characterized by sharp peak (casp) in the field distribution. The third region $(u_{1}^{\prime},u_{2}^{\prime})$, denoted by $\mathcal{U}_{0}$, acts as a separator between the two branes. Here, the field and its derivatives are negligible, and hence this part of the field action $S_{0}\simeq 0$ with appropriate accuracy. The precise definition of the boundaries between the regions $\mathcal{U}_{1}$, $\mathcal{U}_{0}$, and $\mathcal{U}_{2}$ is not necessary.
The first step in application of the Method of Stationary Phase is to find classical function delivering an extrema of action. Assume that this function has two sharp maxima $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ and is (almost) zero between them region $\mathcal{U}_{0}$, like in Fig.2. Then, the action can be segmented into three components based on integration over the regions $\mathcal{U}_{1}$, $\mathcal{U}_{0}$, and $\mathcal{U}_{2}$ Taking all this into consideration, the transition amplitude (22)
$A=N\int D\zeta_{1}\exp{(iS_{1}[\zeta_{1}]})\int D\zeta_{0}\exp{(iS_{0}[\zeta_{0}]})\int D\zeta_{2}\exp{(iS_{2}[\zeta_{2}]})$ (25)
$D\zeta_{k}=\prod_{x,y,u\in\mathcal{U}_{k}}d\zeta_{k}(x,y),\quad k=0,1,2.$ (26)
Measure $D\zeta_{0}$ includes the separating points $\zeta_{1}^{\prime}$ and $\zeta_{2}^{\prime}$. Now we can apply the Method of Stationary Phase for three integral in the same manner as it has been done in (4.2) to obtain
$A=N^{\prime}\int\prod_{x}d\Upsilon_{1}(x)\exp{\left(iS[\Upsilon_{1}(x)Y_{1,cl}(y);\lambda_{in}]\right)}\int\prod_{x}d\Upsilon_{2}(x)\exp{\left(iS[\Upsilon_{2}(x)Y_{2,cl}(y);\lambda_{in}]\right)}$ (27)
The function $\Upsilon(x)$ is denoted as $\Upsilon_{1}(x)$ in the region $\mathcal{U}_{1}$ and similarly for other regions. The second of the three integrals, $S_{0}$, is essentially zero owing to the negligible field value $Y_{cl}(y)$ and its derivatives between the branes (region $\mathcal{U}_{0}$). Knowledge of the classical functions $Y_{1,2,cl}(y)$ allows one to perform the integration of the actions $S_{1}$ and $S_{2}$ over the extra coordinates $y$ which leads to the amplitude
$A=N^{\prime}\int\prod_{x}d\Upsilon_{1}(x)\exp{\left(iS^{(4)}[\Upsilon_{1}(x);\lambda_{eff,1}]\right)}\int\prod_{x}d\Upsilon_{2}(x)\exp{\left(iS^{(4)}[\Upsilon_{2}(x);\lambda_{eff,2}]\right)}$ (28)
for the fields $\zeta_{1},\zeta_{2}$ acting in the disconnected areas $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$. The normalization constant is changed as
$N^{\prime}=N\int D\zeta_{0}.$
The integration here is performed over the spatial domain $\mathcal{U}_{0}$, where $S_{0}=0$.
Formally, transition amplitude (28) describes two independent fields $\Upsilon_{1}$ and $\Upsilon_{2}$ acting in 4-dim. An essential fact is the difference in the parameter values of the fields, $\lambda_{eff,1}\neq\lambda_{eff,2}$, located near different branes. Indeed, the integration over the extra dimensional coordinates $y$ leads to a transformation of the initial parameter values $\lambda_{in}$ in (27) into the effective parameters $\lambda_{eff,1}$ for the fields on brane-1 and $\lambda_{eff,2}$ for brane-2.
The fields in the different branes do not feel each other. However, there are waves that could transfer energy between the branes. At the low energies, the wavelengths of external excitations are much larger compared to the extra space scale and hence cannot be excited.
4.3 The effective parameters of a field are brane-dependent.
The existence of a region characterized by an extremely small amplitude of a field between the branes as in Fig.2, leads the bifurcation of the field into two effective fields with distinct parameters, which has profound implications. This subject is examined in detail in the previous subsection. It is shown therein that a field situated on distinct branes can be treated as two independent fields, each with its own set of the physical parameters.
The explanation on the ’’physical level’’ attracts the analogy with a scalar field behavior within the double-well asymmetric potential. When the energy of the system tends to zero, the field is concentrated near the potential minimums. In this case, the small-amplitude fluctuations around each of the minima are observed as free scalar particles with specific masses depending on the shapes of the minima.
A maximum of potential is responsible for the exponential attenuation of the field amplitude between the two minima. The overlapping integral of the fluctuations near distinct minima is small, so that the fluctuations around both minima do not interact with each other and can be considered as independent fields.
The same is true for fields concentrated near branes. The field near one brane does not ’’feel’’ the field located on the other brane due to the smallness of the overlapping integral. This is true for the zero mode $y_{cl}$. The KK mode excitations are forbidden if their typical wavelength $l\gg l_{extra}$,  ($l_{extra}$ is the distance between the branes) is much smaller than the electroweak scale.
The field distributions near the two branes can be independently approximated by an appropriate sets of orthonormalized functions specific for each brane. These could be the sets built on the basis of harmonic oscillator, or the Coulomb set of wave functions for the hydrogen atom, for example. The general form of the wave function for the n-th energy level of a harmonic oscillator is as follows
$\phi_{n}(u)=N_{n}H_{n}(\alpha u)e^{\alpha^{2}u^{2}/2}$ (29)
where $N_{n}$ is the normalization factor, $H_{n}$ is the Hermite polynomial of the n-th order, $\alpha$ is an arbitrary parameter. In the following, we will assume that all distributions are described by the set of functions $\phi_{n}(u-u_{1})$ if they are located near brane-1, and functions $\phi_{n}(u_{2}-u)$ if they are located near brane-2.
It is assumed that the observers are located at the brane-1. In this case, it is appropriate to choose the field distribution as a sum over the orthonormal basis (29)
$Y_{f}^{(1)}(u)=\sum_{n}a_{n}\phi_{n}(u-u_{1})$ (30)
which reflects the fact of the field localization around $u=u_{1}$ (brane-1). A similar expression can be written for the fermion field located on brane-2 at $u=u_{2}$,
$Y_{f}^{(2)}(u)=\sum_{n}b_{n}\phi_{n}(u_{2}-u).$ (31)
These functions quickly tend to zero if we limit ourselves by several terms in the sums. Thus, the distribution over the whole interval $(u_{1},u_{2})$ can be approximated with good accuracy in the form
$\Psi(x,u)=\Psi^{(1)}(x,u)+\Psi^{(2)}(x,u)=\psi_{1}(x)Y^{(1)}(u)+\psi_{2}(x)Y^{(2)}(u)$ (32)
The first term has a sharp maximum at the brane-1, while the second term has a sharp maximum at the brane-2. This approximation is valid for all sorts of fields, provided they are localized on the different branes. Substitution (32) into the action (17) leads to the effective action for the two free fermions $\Psi^{(1,2)}$. For example, the interference term $\bar{\Psi^{(1)}}iD_{A}\Gamma^{A}\Psi^{(2)}$ is proportional to the overlapping integral
[…]
an hence is negligibly small. The main result of this section is the following: the fermions concentrated around different branes do not influence each other and can be considered as two different sorts of fermions.
4.4 Electromagnetic Field and Gravity between branes
In this subsection, we use the results obtained in [Cui:2020fiz, Popov:2024nax] to show that the electromagnetic field and the gravitational excitations are distributed uniformly between the branes.
We start with the action of the abelian gauge field and the standard form of its D-dimensional action, see e.g. [Dubovsky:2000av]
$S_{gauge}=-\frac{\text{g}_{D}}{4}\int d^{D}X\sqrt{|g_{\text{D}}|}g^{PM}g^{QN}F_{PQ}F_{MN},$
$F_{QN}=\partial_{Q}A_{N}-\partial_{P}A_{Q}.$ (34)
Here $\text{g}_{D}$ is the D-dimensional coupling constant. The classical equations
$\partial_{P}(\sqrt{|g_{D}|}g^{PQ}g^{MN}F_{QN})=0$ (35)
are used to find the field distribution along the extra coordinate $u$. Keeping in mind expression (15) in the form
$A^{N}(x,u)=a^{N}(x)Y_{em}(u)$ (36)
equations (35) can be rewritten as follows ($n=2$)
$M=\mu:\quad a^{\mu}\left[Y_{em}^{\prime\prime}+\left(2\gamma^{\prime}+\frac{r^{\prime}}{r}\right)Y_{em}^{\prime}\right]=0,$
$M=\theta:\quad a^{\theta}\left[Y_{em}^{\prime\prime}+\left(4\gamma^{\prime}-\frac{r^{\prime}}{r}\right)Y_{em}^{\prime}\right]=0.$ (37)
Here, only nontrivial equations are written. Remind that the conditions $\partial_{\mu}\cdot=0,\,\mu=0,1,2,3$ are used. The latter holds at the low energies where 4D wavelengths are much larger than the extra space scale. Evidently, $Y_{em}(u)=C_{em}=const$ is the unique solution to system (4.4) if $a^{\mu},\,a^{\theta}\neq 0$ which means that the electromagnetic field is uniformly distributed between the branes. Charges located on one brane can emit the photons which are absorbed by charges on the other brane, see discussion in section 5.
Let us now discuss the gravitational perturbation over the extra coordinate $u$. The metric containing the tensor gravitational perturbations is of the form
$ds^{2}=e^{2\gamma(u)}(\eta_{\mu\nu}+h_{\mu\nu})dx^{\mu}dx^{\nu}-du^{2}-r(u)^{2}d\phi^{2}$ (38)
$\eta_{\mu\nu}=diag(1,-1,-1,-1),\quad h_{\mu\nu}\ll\eta_{\mu\nu}$ where the perturbation $h_{\mu\nu}$ represents the tensor mode obeying the transverse-traceless gauge conditions
$\eta_{\mu\nu}h^{\mu\nu}=0,\quad\partial_{\mu}h_{\mu}^{\nu}=0.$
During this paper, we are interested in the main term of the decomposition (15) and represent the perturbations in the form
$h_{\mu\nu}(x,u)=H^{(0)}_{\mu\nu}(x)Y_{0}(u)$ (39)
Our aim is to find the distributions over the extra dimensions so that we put $H^{(0)}_{\mu\nu}(x)=const$ (it is assumed that the $x-$derivatives are small, or $\partial_{\mu}\cdot=0$ ). After some algebra, see [Cui:2020fiz], equations for the linear tensor perturbations are reduced to a non-trivial equation
$\left[\left(4\gamma^{\prime}+\frac{r^{\prime}}{r}\right)\partial_{u}+\partial_{u}^{2}+\frac{\partial_{u}f_{R}}{f_{R}}\partial_{u}\right]Y_{0}(u)=0.$ (40)
for the tensor mode (39). One of the possible solutions of the equation (40) is $Y_{0}=const$, as it should be for the zero mode. This means that the gravitational field is uniformly distributed between the two branes. Matter on one brane experiences the gravitational forces produced by matter on the other brane.
5 Manifestation of the second brane
In this section, we will briefly examine the possible observational implications of the hypothetical sandwich model of extra dimensions discussed in this paper. In the previous section, we have shown that
- both branes are filled with similar types of fields;
- the photons and the gravitational waves in 4D are distributed uniformly over the extra dimensions;
- The hierarchy problem might be solved exclusively within a single brane framework, implying that the formation of complex astrophysical structures — such as stars — could be restricted to brane-1.
5.1 Photon interaction
The subsection 2.2 deals with the challenges of interpreting these particles as dark matter candidates. The most obvious property of dark matter is its extremely weak interaction with the electromagnetic field. Here we will answer two questions: are the charges of the electrons on different branes equal, and can such electrons form a bound state? So, we have to consider a D-dim extension of the SM action for the electron-photon interaction
[…]
Here $\alpha_{D}$ is a coupling constant in D dimensions and $A_{N}$ is the gauge field. It is shown above that the gauge field is uniformly distributed over the extra dimensions, i.e. $A_{N}(x,y)=C_{em}a_{N}(x)$. Action (41) can be reduced to the 4-dim action:
$S_{int}\simeq\alpha_{D}{v}_{\text{n}-1}\int\sqrt{|g_{4}|}d^{4}x\int_{u_{1}}^{u_{2}}du\sqrt{|g_{n}|}\left(\bar{\Psi}_{1}\Gamma^{N}A_{N}\Psi_{1}+\bar{\Psi}_{2}\Gamma^{N}A_{N}\Psi_{2}+\bar{\Psi}_{1}\Gamma^{N}A_{N}\Psi_{2}+h.c.\right)$ (42)
where $\Psi=\Psi_{1}+\Psi_{2}$ and the functions $\Psi_{1}$, $\Psi_{2}$ are located on the brane-1 and brane-2 correspondingly, according to discussion in subsection 4.2. The next step is to substitute the decompositions $\Psi_{1,2}=\psi_{1,2}(x)\cdot Y_{f}^{(1,2)}(u)$ into action (42), to obtain
$S_{int}=\int\sqrt{|g_{4}|}d^{4}x\left(\alpha\bar{\psi}_{1}\gamma^{\mu}a_{\mu}\psi_{1}+\alpha\bar{\psi}_{2}\gamma^{\mu}a_{\mu}\psi_{2}+\alpha_{12}\bar{\psi}_{1}\gamma^{\mu}a_{\mu}\psi_{2}+h.c.\right)$ (43)
where expressions (19) and (32) have been kept in mind. We also used the fact that the extra space distribution of the fermions is included into the normalization of the wave functions $\psi_{1,2}$ and $\alpha=C^{2}_{em}\alpha_{D}$. The observable coupling constant $\alpha$ is the same on both branes. The parameter $\alpha_{12}$ is proportional to the overlapping integral (33) and hence, is extremely small.
Action (43) contains the vertices describing the interaction of photons with ’’normal’’ electrons $\psi_{1}$ located at the brane-1 (first term) and ’’heavy’’ electrons $\psi_{2}$ on the brane-2 (second term). A photon emitted from the brane-1 can be absorbed by fermions located on brane-2, and vice versa. We will use this fact below.
The third term represents the vertex containing both types of electrons in conjunction with the electromagnetic field. The transition of the heavy electron to another brane, accompanied by photon emission, is associated with this vertex. The latter could also lead to the bound states containing the two electrons on different branes. Both of these effects are negligibly small due to the smallness of the coupling constant $\alpha_{12}$.
5.2 The Higgs mechanism on different branes
As was mentioned in subsection 2.2, the brane other than that settled by observers are most likely filed by heavy partners of the SM particles. According to the Standard Model, the Higgs field is responsible for the fermion masses. Therefore, it is of interest to study the extra dimensional distribution of the Higgs field and its interaction with the fermions.
The distribution of the Higgs field
$H(x,y)=h(x)Y_{H}(u),\quad h(x)=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{h}+\rho(x)\end{pmatrix},\quad v_{h}\gg\rho(x)$ (44)
in the form (15) over the extra coordinate $u$ is governed by the scalar function $Y_{H}(u)$ which is the solution to the classical equation
$-\left[\partial^{2}_{u}+\left(4\gamma^{\prime}+\dfrac{r^{\prime}}{r}\right)\partial_{u}\right]Y_{H}(u)=\nu Y_{H}(u)-\lambda\,Y_{H}^{3}(u),$ (45)
As was shown in [Bronnikov:2023lej], the field distribution over the extra dimensions is specific to each causally separated 3D region below the horizon. Consequently, the field distributions are unique for each universe generated during the de Sitter stage. One of the solutions is shown in Fig. 3.
Figure 3: The Higgs field distribution over the internal coordinate $u$ is specific to each brane. The background metric functions are represented in Fig.1. Additional conditions are $Y(1.5)=0,\,Y^{\prime}(1.5)=10^{-2}$.
The expressions for the Higgs mass and the coupling constant as functionals of the function $Y_{H}(u)$ were obtained in [Bronnikov:2023lej].
[…]
Here $\nu,\,\Lambda$ are the initial physical parameters of the Higgs field in D-dim, ${v}_{\text{n}-1}$ is the volume of the $(\text{n}-1)$-dimensional sphere. This is the specific example of the renormalization of the initial physical parameter values $\lambda_{in}\to\lambda_{eff}$ introduced in subsection 4.2, where $\lambda_{in}=\nu,\,\Lambda$ and $\lambda_{eff}=m_{h},\,\lambda_{h}$.
The Higgs distribution $H$ can be split as $H(x,y)=H_{1}(x,u)+H_{2}(x,u)$, see general expression (32). Recall that the distributions $H_{1}(x,u)$ and $H_{2}(x,u)$ act as independent fields as is discussed in 4.2 and 4.3. The function $H_{1}(x,u)$ decreases rapidly far from the endpoint $u_{1}$ so that the observer on the brane-1 perceives the Higgs field as $H(x,y)=H_{1}(x,u)=h_{1}(x)Y^{(1)}_{H}(u)$. The function $H_{2}(x,u)$ quickly decreases far from the endpoint $u_{2}$ and a possible observer on the brane-2 perceives the Higgs field as $H(x,y)=H_{2}(x,y)=h_{2}(x)Y^{(2)}_{H}(u)$. Therefore, the 4D parameters of the Higgs field $m_{h},\lambda_{h}$ are functionals of the distribution $Y^{(1)}_{H}(y)$ for the brane-1 observer and $Y^{(2)}_{H}(y)$ for the brane-2 possible observer. Hence, they are different on both branes. The same is true for the Higgs vacuum average
$v_{h}=\frac{m_{h}}{\sqrt{2\lambda_{h}}}.$ (48)
It was declared from the beginning that the model is free from initially small physical parameter values. This means that the integrals as in (46) and (47) are of the order of unity if no special effort was made. The fine-tuning should be done in the expression (46) to obtain the observable Higgs mass and its vacuum average. As shown in the paper [Bronnikov:2023lej], the vacuum average can be adjusted to the known value $v_{h}\sim 10^{-17}m_{Planck}$ by an appropriate selection of the classical distribution $Y_{H}(u)$ over the extra dimensions. In the framework of the two-brane model discussed here, this procedure can be done for the brane-1 where the observer is assumed. This means that $Y_{H}(u)$ should be substituted by $Y^{(1)}_{H}(u)$ in formulas (46) and (47). The Higgs vacuum average on the brane-2 remains uncertain.
In the Standard Model, the fermion masses are proportional to the Higgs vacuum average $v_{h}$. The fermion-Higgs vertex in $D$ dimensions has the form
$S_{He}=\int d^{D}X\sqrt{g_{D}}f_{E}\bar{\Psi}_{L}H\Psi_{R}$ (49)
where ${\Psi}_{L}$ is a left doublet of the SU(2) group, $\Psi_{R}$ is a right singlet and $H$ is the Higgs field. Below we show that the fermion masses and their coupling constant with the Higgs are different on both branes in 4D.
The decompositions (32) in the form (15)
$\Psi_{L}^{(1,2)}=\psi_{L}(x)Y^{(1,2)}_{L}(u),$ (50)
$\Psi_{R}^{(1,2)}=\psi_{R}(x)Y^{(1,2)}_{R}(u)$ (51)
of the fermions in D-dim on brane-1 (marked as $Y^{(1)}_{L,R}$) and brane-2 (marked as $Y^{(2)}_{L,R}$) leads to the following action
$S^{(1,2)}_{He}=\int d^{4}x\sqrt{g_{4}}f_{e}^{(1,2)}\bar{\psi}_{L}(x)h(x)\psi_{R}(x)$ (52)
after their substitution in (49). All functions acting in the 4D space are assumed to be normalized to obtain the standard form of the kinetic terms. The value of the 4D coupling constant
[…]
depends on which brane the fermion is located because the fermion distributions reveal a sharp peaks: $Y^{(1)}_{L,R}(u)$ on the brane-1 and $Y^{(2)}_{L,R}(u)$ on the brane-2, see Fig.2. The metric functions $\gamma(u),r(u)$ and the distribution $Y_{H}(u)$ are unique near each brane, so the coupling constant values $f^{(1,2)}_{e}$ are different there.
If we live on brane-1, the SM in 4 dimensions and the parameter $f^{(1)}_{e}$ must have the observable value $f^{(1)}_{e}=m_{e}/v^{(1)}_{h}\sim 10^{-6}$, see [Workman:2022ynf]. This value is very small and hence needs to be fine-tuned in addition to the vacuum average $v_{h}$. This can be done by selecting appropriate functions $Y^{(1)}_{L,R}(u)$ in (53).
In this research, we concentrate on the fields located on the brane-2, where the vacuum average $v^{(2)}_{h}$, the coupling constant $f^{(2)}_{e}$ and hence, the electron mass $m_{2}\sim f^{(2)}_{e}v^{(2)}_{h}$ remain uncertain varying in a wide range due to the absence of the fine-tuning.
5.3 Ultra-high-energy cosmic rays
Ultra-high-energy cosmic rays (UHECR) of the energy $10^{11}$eV [ParticleDataGroup:2024cfk] are one of the most intriguing mysteries of modern astrophysics. So far, only several protons of such energy were definitely registered. Despite numerous attempts to explain this phenomenon, its origin remains largely unknown. The discussion on the origin of the high-energy photons of PeV energy can be found in [2011APh....35...28B].
The two-brane model discussed here provides us with a possible explanation of this phenomenon based on the heavy electron-positron annihilation into a pair of quarks. The part of action (43) contains three vertexes. The first term contains two fermion settled on brane-1 interacting with an electromagnetic field. The second one contains two fermions settled on the neighboring brane-2 coupled with the same electromagnetic field. So, the process describing the annihilation fermions on brane-2 into photon with the following creation of two other fermions on the brane-1 can be realized. The last term in (43) is proportional to the extremely small coupling constant $\alpha_{12}$ and is hardly significant. The energy $\varepsilon$ of the emitted quarks on the brane-1 is of the order of the heavy electron mass $m_{2}$
$\varepsilon\sim m_{2}$ (54)
provided that they are non-relativistic. This allows us to estimate the mass of the electrons settled on the brane-2 if we suppose that UHECR are the result of the heavy electron-positron annihilation,
$m_{2}\sim 10^{11}\text{GeV}.$ (55)
This value is consistent with the lower boundary limits (4). The small number of such events is due to the extremely small cross-section, as previously discussed.
A similar mechanism used to explain high-energy photons encounters difficulties. The highest observed photon energy is about $10^{6}$ GeV, and the mass of a heavy electron must be of this order of the magnitude. However, this value is below the limit marked in (4), so some efforts should be applied to relate this effect to the dark matter.
6 Conclusion
Two-brane extra dimensions play an important role in theoretical research starting from the Randall-Sundrum model. In this research, we study the distribution of fields using the two-brane model [Popov:2024nax]. This model has revealed its ability to explain the electroweak hierarchy problem, the restoration of the Starobinsky inflationary model [Starobinsky:1980te], and the smallness of the cosmological constant [Bronnikov:2023lej].
Here we focus on the fields distribution over the extra dimensions. The extra-dimensional regions where fermions are localized are referred to as (thick) branes. It is shown that fermions split into two independent effective fields, localized on adjacent branes. This bifurcation process accompanies the evolution of the Universe from energy levels above the inflationary scale to the present-day state. The conditions under which the fermions and the Higgs field are independent at neighboring branes are attentively discussed in the subsection 4.2. At the same time,the gauge fields are uniformly distributed in the extra dimensions, preserving the charge universality in 4D physics.
The emergence of complex structures requires a hierarchy of energy scales and associated fine-tuning [2007unmu.book..231D]. In a brane-world model, the delicate process of parameter selection for this fine-tuning can be accomplished on a single brane, the one inhabited by the observers (brane-1). Consequently, a secondary brane (brane-2) would lack this fine-tuning, precluding the formation of stellar-like structures. This means that the masses of the particles on the other brane should have ’’natural’’ values, i.e. be of the order of the energy scale at which our Universe was formed, i.e. extremely massive. The Higgs mechanism of mass creation on both branes is discussed in subsection 5.2.
The stable matter located on the brane-2 most plausibly consists of the charged particles such as 3 generations of leptons and the protons as well as neutral particles like the neutrino and the atoms of hydrogen. Neutral particles are obvious candidates for the role of dark matter. However, the heavy-charged leptons are also promising candidates. The extremely high mass of such leptons on brane-2 is the key factor for the significant reduction of the Compton scattering cross section. This allows the heavy leptons to be considered as a component of dark matter.
It is shown that the fermions located on the different branes can exchange via photons and gravitons and thus be a source of ultra-high energy particles.
The Sandwich model of extra dimensions discussed here seems to be a promising framework that deserves further development. It could also serve as a basis for explaining phenomena such as the positron anomaly [Belotsky:2018vyt] and the smallness of the neutrino mass.
Acknowledgements
The author is grateful to K. Belotsky and R. Konoplich for fruitful discussions. The work was partly funded by the Ministry of Science and Higher Education of the Russian Federation, Project "New Phenomena in Particle Physics and the Early Universe" FSWU-2023-0073. The research presented in Sect.3,4 was also partially carried out in accordance with the Kazan Federal University Strategic Academic Leadership Program.
[…]
[title=References, heading=bibintoc]